OpenStudy (logan13):

What is the solution to the system of equations represented by these two lines? A. (3, 1) B. (3, 0) C. (1, 3) D. (0, 6)

OpenStudy (logan13):

@dan815

OpenStudy (logan13):

@pooja195

OpenStudy (logan13):

@pooja195 plz help

OpenStudy (logan13):

@ShadowLegendX

OpenStudy (shadowlegendx):

You need to find the slopes and y intercepts of both lines

OpenStudy (logan13):

so c

OpenStudy (shadowlegendx):

They should come out into a format of... y = mx + b with m being the slope and b being the y intercept

OpenStudy (shadowlegendx):

Well what did you get for the two lines?

OpenStudy (shadowlegendx):

y = -2x + 7

OpenStudy (shadowlegendx):

y = 1/3x

OpenStudy (shadowlegendx):

Thatʻs what I got, what did you get

OpenStudy (logan13):

so c right

OpenStudy (shadowlegendx):

If itʻs c... (1,3) youʻre saying x = 1 and y = 3 3 = 1/3(1) That is not a correct equations

OpenStudy (shadowlegendx):

I know the answer, but it isnʻt C

OpenStudy (logan13):

@ShadowLegendX A.

OpenStudy (anonymous):

i think that it is a